OpenStudy (anonymous):

What is the solution to the equation x + 2.15 = –3.2?

alones (alones):

First of all subtract 2.15 from both sides. x+2.15−2.15=−3.2−2.15 x=−5.35

OpenStudy (scooby6363):

Subtract 2.15 from both sides. When you substract a negative number with a positive you are essentially "adding" both numbers together. So -3.2 - 2.15 = -5.35 This would be the solution as now x = -5.35 When plugging in -5.35 for x you get: -5.35 + 2.15 = -3.2

OpenStudy (anonymous):

thanks

OpenStudy (gabbyalicorn):

YOUR PROBLEM: x+2.15=−3.2 SUBTRACT 2.15 FROM BOTH SIDES: x+2.15−2.15=−3.2−2.15 x=−5.35 ANSWER: x=−5.35